OpenStudy (sonyalee77):

The ratio 3:4 is the same as the ratio of 4:3. A. True B. False

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

3:4 is \(\frac{3}{4}\) and 4:3 is \(\frac{4}{3}\).. is their any difference?

OpenStudy (anonymous):

@Sonyalee77 I am expecting a reply from your side..

OpenStudy (sonyalee77):

so sorry I had to do something for someone but to me in this problem the only difference is that the 3 and the 4 are in different places

OpenStudy (anonymous):

yes but their impact is also different.. \[\frac{3}{4} = 0.75\] \[\frac{4}{3} = 1.33\] Now is their any difference??

OpenStudy (anonymous):

Are they equal??

OpenStudy (sonyalee77):

oh I should have thouht of that yes there is a difference THANK YOU SOO MUCH!!!

OpenStudy (anonymous):

So what should be your answer??

OpenStudy (sonyalee77):

false

OpenStudy (anonymous):

True...

OpenStudy (anonymous):

I mean you are true.. Ha ha ha ha, yep you are right...

OpenStudy (sonyalee77):

lol you scared me for minute

OpenStudy (anonymous):

Ha ha ha...